OpenStudy (breanna.rae):

What is the orthocenter of the triangle with two altitudes given by the lines x = -1 and y = x+1

OpenStudy (breanna.rae):

@MaxwellFish

OpenStudy (maxwellfish):

1-Find the equations of 2 segments of the triangle (for our example we will find the equations for AB, and BC) 2-Once you have the equations from step #1, you can find the slope of the corresponding perpendicular lines. 3-You will use the slopes you have found from step #2, and the corresponding opposite vertex to find the equations of the 2 lines. 4-Once you have the equation of the 2 lines from step #3, you can solve the corresponding x and y, which is the coordinates of the orthocenter. 5- Find equations of the line segments AB and BC. To find any line segment, you will need to find the slope of the line and then the corresponding y-intercept. A (3, 1) B(2, 2) C (3, 5) Slope of AB = (1-2)/(3-2) = -1/1 = -1 y = mx + b (substitute m = -1, x = 3, y = 1) 1 = -1(3) + b b = 4 Equation of AB: y = -1x + 4 Slope of BC = (2-5)/(2-3) = -3/-1 = 3 y = mx + b (substitute m = 3, x = 2, y = 2) 2 = 3(2) + b b = -4 Equation of BC: y = 3x - 4 6- Find the slope of the corresponding perpendicular lines Slope of AB = -1 Slope of perpendicular line to AB: -1*m = -1 -> m = 1 Slope of BC = 3 Slope of perpendicular line to BC: 3*m = -1 -> m = -1/3 7- Find the equation of the perpendicular lines Slope of perpendicular line to AB: m = 1 We will use the coordinate of the opposite vertex (point C) to find the equation of the line. y = mx + b (substitute m = 1, x = 3, y = 5) 5 = 1(3) + b b = 2 Equation of perpendicular line to AB: y = 1x + 2 Slope of perpendicular line to BC: m = -1/3 We will use the coordinate of the opposite vertex (point A) to find the equation of the line. y = mx + b (substitute m = -1/3, x = 3, y = 1) 1 = -1/3*(3) + b b = 2 Equation of perpendicular line to AB: y = -1/3x + 2 8- solve 2 perpendicular lines equation 1: y = 1x + 2 equation 2: y = -1/3x + 2 Solving for x and y: 1x + 2 = -1/3x + 2 4/3x = 0 x = 0 y = 1(0) + 2 y = 2 The coordinates are (0, 2). This is the orthocenter.

OpenStudy (maxwellfish):

my teacher gave me that

OpenStudy (maxwellfish):

hope that helps

OpenStudy (breanna.rae):

omg thank you so much!!

OpenStudy (maxwellfish):

No Problem